A population consists of the numbers 1, 2, 3, and 4 with the sample size of 2. Compute the population variance.




1
Expert's answer
2022-04-14T04:36:03-0400

"\mu =\frac{1+2+3+4}{4}=2.5\\\sigma ^2=\frac{\left( 1-2.5 \right) ^2+\left( 2-2.5 \right) ^2+\left( 3-2.5 \right) ^2+\left( 4-2.5 \right) ^2}{4}=1.25"